Jethroe Co. reported a retained earnings balance of $200,000 at December 1, year 4. In June year 5, Jethroe discovered that merchandise costing $50,000 had not been included in inventory in its year 4 financial statements. Jethroe has a 21% tax rate.

What amount should Jethroe report as adjusted beginning retained earnings in its statement of retained earnings at December 31, year 5?

Multiple Choice

  • $150,000

  • $189,500

  • $200,000

  • $239,500
Answer #1

Answer

--Inventory omitted to be recorded will lead to Net Income (and Retained earnings) being understated.

--When inventory is added to books of account, Net Income of previous year will increase. Tax rate is 21% and hence, tax on increased net income because of omitted inventory will be deducted.

--Increase in Retained earnings balance, net of tax = $ 50000 - (50000 x 21%)
= 50000 - 10500
= $ 39,500

--Adjusted beginning retained earnings = 200000 + 39500 = $ 239,500

--Correct Answer Option #4: $ 239,500